y= -3x
y= x-8
Solving linear equations using the substitution method

Answer by mananth(16946)

y= -3x
y= x-8
substitute y in the second equation
-3x=x-8
add -x to both sides
-4x=-8
/-4
x=2
plug x =2 in y=-3x
y=-3*2
y=-6